what is slope of a line that passes throught(1,1)(1,2)
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

slope = (y2-y1)/(x2-x1)
slope = 2-1/1-1
slope = 1/0
slope not defined.
....
m.ananth@hotmail.ca

Answer by stanbon(75887)

what is slope of a line that passes throught(1,1)(1,2)
-------------
slope = change in y/change in x
---
= (2-1)/(1-1)
= 1/0
which is "undefined".
---
If you plot the two points and draw a line
thru the 2 points you will see that you
get a vertical line that contains all of
the points where x = 1.
That is the equationof the "line" and
it has an "undefined" slope.
===========
Cheers,
Stan H.